### 09 Mar 2026 v1

@@ 1 Introduction
Whataboutism is a tactic that people sometimes use to deflect a criticism against violations of social norms. Merriam-Webster dictionary defines whataboutism as “the act or practice of responding to an accusation of wrongdoing by claiming that an offense committed by another is similar or worse.” For example, when a prominent public figure from camp A is reprimanded by a member of camp B for mocking the spouse of a camp-B politician (or for keeping silent when another prominent member of camp A did so), a whatboutism response can take the following form: “Why are you criticizing me? Remember when a prominent member of your party ridiculed the child of a politician from my camp? I don’t recall seeing anybody from your party condemning her.”
Whatboutism is a modern-day instance of an ancient rhetorical device known as tu quoque — which Merriam-Webster dictionary defines as “a retort charging an adversary with being or doing what the adversary criticizes in others.” There is a popular sentiment that whatboutism has become increasingly common in recent decades (e.g., see Dykstra (2020)), and that it has sullied public discourse and eroded the liberal ethos that we can agree over means even when we disagree over ends. In Sucio (2025), the Forbes journalist Peter Sucio wrote: [^1] https://www.forbes.com/sites/petersuciu/2025/04/04/whataboutism-has-made-civil-debate-on-social-media-nearly-impossible/ “No matter the topic – especially those about politics and world affairs – it is impossible to have any civil debate on social media without someone responding by making a counter-accusation. This isn’t new, but it has taken such a deep foothold in American politics in recent years, and it isn’t hard to see why.” Sucio went on to attribute the spread of whataboutism to the rise of social media and political polarization.
We propose a formalization of whataboutism as an equilibrium phenomenon, and use this formalism to explore the driving forces behind the phenomenon and its implications for public discourse. In constructing our model, we are guided by three desiderata. First, context should matter: Insulting a soldier on Memorial Day is more offensive than on normal days, and whataboutism arguments would make use of this distinction. Second, the model should enable criticism to be directed not only at offensive speech, but also at subsequent failures to condemn it from within the offender’s camp. Finally, availability of the whatboutism argument should be endogenous: The scenario referred to by one camp’s whataboutism argument is drawn from past speech acts performed by agents on the other camp, which in turn obeyed the same norms shaped by whataboutism. This means that in our model, the social cost of offensive speech should be a function of agents’ equilibrium beliefs regarding the behavior of others.
@@
A condemnation’s effectiveness depends on whether it is internal or external. Condemnation by a member of one’s own camp is always effective; it automatically inflicts a cost on an agent who performed/supported offensive speech. In contrast, external condemnation can be rebuffed with a whataboutism argument. The argument is based on drawing a play path from the distribution induced by equilibrium strategies, in a state that satisfies three requirements: $(1)$ the state is at least as sensitive as the current one; $(2)$ the initial speech-act choice is faced by an agent from the rival camp; and $(3)$ that agent performs an offensive speech act and meets no internal rebuke along the play path. This definition captures whataboutism arguments, which take the following general form: “In a different context that is at least as sensitive as the present one, someone from your camp performed an offensive speech act and no one from your camp condemned it.” The agent’s cost from external condemnation is defined as the probability that he cannot deflect it with whatboutism. The fact that the cost is determined by the ex-ante equilibrium distribution over play paths is what turns our model into a psychological game.
@@ 2 A Model
There are two rival camps, $1$ and $2$, each consisting of a measure one of agents. We use i-i to denote the rival of camp $i$. There is a set of states $S=S_{1}\cup S_{2}$, where $S_{i}=\left\{s_{i}^{1},...,s_{i}^{n}\right\}$ and $S_{1}\cap S_{2}=\emptyset$. Denote $I(s)=i$ when $s\in S_{i}$. In state $s$, camp $I(s)$ gets an opportunity to engage in offensive speech against the other camp. An offensive speech act in state $s$ gives intrinsic utility $v\sim U[0,g_{m}]$ to a member of camp $I(s)$, and intrinsic disutility drawn from $U[0,b_{m}]$ to a member of camp I(s)-I(s). Assume $1<g_{n}<\cdots<g_{1}<2$ and $1<b_{1}<\cdots<b_{n}<2$. A larger $m$ represents greater sensitivity to offensive speech: It generates lower intrinsic utility for offenders and higher disutility for victims. When $m^{\prime}\geq m$, we sometimes write $s_{i}^{m^{\prime}}\succsim s_{j}^{m}$ for any $i,j$, and say that $s_{i}^{m^{\prime}}$ is at least as sensitive as $s_{j}^{m}$.
@@
Agents from camp I(s)-I(s) follow an exogenous rule (which we endogenize below): Condemning camp $I(s)$ with probability $\lambda b_{s}$, where $\lambda\in(0,\frac{1}{2})$. The only endogenous objects are the strategies of camp-$I(s)$ agents. Since agents from the rival camp I(s)-I(s) are selected at each stage with probability $\frac{1}{2}$ and terminate the game with probability $\lambda b_{s}$, it follows that any profile $\sigma$ of strategies by camp-$I(s)$ agents induces a well-defined probability distribution over finite terminal histories. For every $s\in S$, let $\alpha_{\sigma}(s)>0$ denote the ex-ante probability that the agent who terminates the game in $s$ is from camp I(s)-I(s), under the strategy profile $\sigma$.
@@
  • Finally (this is where the psychological-game component enters), when the agent plays $a=1$ and his immediate successor is from the rival camp I(s)-I(s) and plays $a=0$, the agent’s payoff is defined relative to a fixed strategy profile $\sigma$. Specifically, his payoff is $v$ if he can mount a “rebuttal”, and v1v-1 if he cannot. A rebuttal consists of a choice of a state $s^{\prime}\succsim s$ for which $I(s^{\prime})=-I(s)$, and a terminal history in that state — drawn randomly from the distribution over terminal histories induced by $\sigma$ — such that according to $\sigma,$ the stage-$1$ agent in that history plays $a=1$, and all subsequent agents from camp $I(s^{\prime})$ play $a=1$ in that terminal history.
@@
The interpretation of the psychological-game component is as follows. When an agent from camp I(s)-I(s) criticizes an agent from camp $I(s)$ for performing/supporting an offensive speech act, the latter agent can employ a “whataboutism” counterargument: In circumstances that are at least as sensitive as $s$, an agent from the critic’s own camp engaged in offensive speech and faced no subsequent internal rebuke. The psychological-game component arises because the rebuttal is based on a random draw from the distribution over play paths induced by the strategy profile $\sigma$. The rebutting agent needs to retrieve a concrete scenario that would arm him with the whatboutism argument. In this scenario, the state needs to be at least as sensitive as $s$; an agent from the rival camp I(s)-I(s) needs to perform an offensive speech act; and in the ensuing play path, there needs to be no internal condemnation, such that the agent who terminates the game is from camp $I(s)$.
Our modeling of the agent’s retrieval process is based on two desiderata: We want to assume that the agent has some control over the retrieval, but also that success is uncertain. Consequently, our model mixes elements of “directed” and “random” search, in analogy to the macro-theory literature (see Wright et al. (2021)). The agent deliberately targets a specific state, but then he — or someone from his own camp who argues on his behalf — randomly retrieves from memory a concrete state-specific scenario. Thus, when the agent plays $a=1$, he does not know for sure whether he will be able to successfully use whatboutism to rebuff out-group condemnation, but his choice of which state to target is optimal in this regard.
@@
Why do we interpret $a$ as a speech act, rather than as a physical action? Imagine that the whatboutism rebuttal were infeasible. In this case, we would be able to interpret $a=1$ as a physical action that benefits one’s own camp and harms the rival camp. An agent who takes this action incurs a cost if his immediate successor plays $a=0$. This is a fairly standard “network externality”: The agent’s benefit from $a=1$ depends on whether other agents take this action, too.
@@ 3 A Benchmark
There is a unique Subgame Perfect Equilibrium in the benchmark model. In each state $s$, each camp-$I(s)$ agent plays $a=1$ if and only if his intrinsic payoff $v$ is above the cutoff
@@
Suppose $v^{k}=0$. This means that camp-$I(s)$ agents with $v<\frac{1}{2}\lambda b$ play $a=1$. Such agents earn a negative payoff, regardless of $v^{k+1}$, a contradiction. Now suppose $v^{k}=g$. This means that camp-$I(s)$ agents with $v>1$ play $a=0$, even though playing $a=1$ would generate a strictly positive payoff, a contradiction. $\square$
@@
The equilibrium fraction of camp-$I(s)$ agents who do not perform/support offensive in state $s$ is
@@
This fraction is always in $(0,1)$. The fact that $x_{s}>0$ for every $s$ means that norms of public discourse never break down entirely — i.e., there is always a positive fraction of agents who refrain from offensive speech or criticize it internally. Note that $x(s)$ is increasing in the parameter $b_{s}$ that determines the disutility from offensive speech that members of camp I(s)-I(s) incur. It is also increasing in the parameter $\lambda$ that affects camp-$I(s)$ agents’ exposure to external criticism. Finally, it is decreasing in the parameter $g_{s}$ that determines the intrinsic utility that camp-$I(s)$ agents derive from verbally attacking the rival camp.
@@ 4 Analysis: Whataboutism
This section characterizes equilibrium behavior in the main model. As a preliminary step, we derive an expression for agents’ psychological expected payoffs. Fix a strategy profile $\sigma$ and consider a state $s$. Consider a camp-$I(s)$ agent with valuation $v$, who is selected to move in some stage $k$. If the agent chooses $a=0$ and thus terminates the game, his payoff is by definition 0. Suppose now that the agent chooses $a=1$. Let $\beta_{\sigma}$ denote the probability that the stage-$(k+1)$ agent will choose $a=0$, given the history, conditional on being from the same camp $I(s)$. Recall that $\alpha_{\sigma}(s^{\prime})$ is the ex-ante probability (induced by $\sigma$) that the agent who terminates the game in $s^{\prime}$ is from camp $I(s^{\prime})$. This is the probability that a randomly sampled state-$s^{\prime}$ play path will support a whataboutism rebuttal by a member of camp $I(s^{\prime})$. Then, the stage-$k$ agent’s psychological expected payoff given $\sigma$ from playing $a=1$ is
@@
The first term in this expression is the agent’s intrinsic payoff from performing/supporting the speech act. The second term represents the probability that the agent faces condemnation from within his own camp, which by assumption is impossible to deflect. The third term represents the probability that the agent faces external condemnation which he cannot rebut with a whatboutism argument. Note that the minimum operator is due to the agent’s ability to select among the eligible states (i.e., those that are at least as sensitive as $s$) to maximize the probability of being able to say: “But what about state $s^{\prime}$? An agent from your camp performed an offensive speech act in that state, which is at least as sensitive as $s$, and I didn’t see anyone from your camp criticizing him.”
@@
Let $S=\{s_{1},s_{2}\}$ where $s_{1}\sim s_{2}$, such that both states are associated with the same $(g,b)$. We focus on stationary and symmetric PSPE — i.e., there is a cutoff $v^{\ast}$ such that at every $s$, a camp-$I(s)$ agent with valuation $v$ chooses $a=1$ at any history, if and only if $v\geq v^{\ast}$. In such an equilibrium, $\beta$ (the probability that a camp-$I(s)$ agent plays $a=0$) is given by $v^{\ast}/g$, independently of the history. Therefore,
@@
The fraction of camp-$I(s)$ agents who play $a=0$ in state $s$ at any stage is $x^{\ast}=v^{\ast}/g$. Note that since $g>1$ and $b\lambda<1$, this fraction $x^{\ast}$ is equal to the benchmark level (4) multiplied by a “discount factor” $b\lambda-2g+2<1$. This factor captures an equilibrium effect of whataboutism, which raises the fraction of agents who perform/support offensive speech acts. When $b\lambda-2g+2<0$ there is no interior solution, and $x^{\ast}=0$ in the unique equilibrium — i.e., all norms that constrain public expression break down.
@@ 4.1 The Main Result
For every $m^{\ast}\in\{M,...,n+1\}$, there is a PSPE such that: (i) for every $m<m^{\ast}$, camp-$I(s_{i}^{m})$ agents always play $a=1$ in state $s_{i}^{m}$; (ii) for every $m\geq m^{\ast}$, a camp-$I(s_{i}^{m})$ agent with valuation $v$ plays $a=1$ if and only if $v\geq v_{m}^{\ast}=c_{m}\theta_{m}g_{m}$. There are no other equilibria.
@@
By essentially the same arguments as in Step 1 of the proof of Proposition 1, for very state $s$ and every stage $k$, there is a cutoff $v_{s}^{k}$ such that a camp-$I(s)$ agent with valuation $v$ who moves in stage $k$ plays $a=1$ if and only if $v\geq v_{s}^{k}$.
Given a PSPE $\sigma,$ $\mu_{s}\left(\sigma\right):=\min_{s^{\prime}\succsim s\mid I(s^{\prime})=-I(s)}(1-\alpha_{\sigma}(s^{\prime}))$ is the probability that external condemnation of camp $I(s)$ in state $s$ cannot be rebuffed by whatboutism. Then, in state $s$, a camp-$I(s)$ agent with valuation $v$ who moves in stage $k$ earns a payoff of
@@
By Step 1, in equilibrium, every state $s$ is associated with a stationary cutoff $v_{s}^{\ast}$ that governs camp-$I(s)$ agents’ behavior. Therefore, for every state $s^{\prime}$ where $I(s^{\prime})\neq I(s)$, we obtain essentially the same expression as (6), repeated here for convenience:
@@
It is now straightforward to confirm that the sequences $(x_{m})$ and $(y_{m})$ described by Step 4 solve (11). Therefore, Step 4 fully characterizes the set of PSPE. $\blacksquare$
Proposition 2 means that any PSPE is characterized by a “sensitivity threshold” $m^{\ast}\geq M$. In state $s_{i}^{m},$ $m\geq m^{\ast}$, the probability that a camp-$i$ agent refrains from offensive speech (and condemns his comrades when they do not) is $x_{m}=v_{m}^{\ast}/g_{m}$, which is equal to the corresponding probability in the benchmark ($c_{m}$) multiplied by the “discount factor” $\theta_{m}.$ Since both $c_{m}$ and $\theta_{m}$ increase with $m,$ the probability of condemnation increases with the sensitivity of the state. The term $\theta_{m}$ can be interpreted as the “whataboutism effect”, which lowers the likelihood that agents condemn members of their camp for performing/supporting offensive speech. The whataboutism effect can be dramatic and lead to a “breakdown of norms” in the sense that in if $\theta_{m}<0$ for some state $s_{i}^{m},$ then in all weakly less sensitive states, all agents (regardless of their camp) perform/support offensive speech.
@@
The PSPE are ranked by their amount of offensive speech: The higher $m^{\ast}$, the higher that amount. In particular, there is always an equilibrium with complete breakdown of norms, where $m^{\ast}=n+1$ such that agents always engage in (and never condemn) offensive speech. To see why, note that the equilibrium probability that in state $s$ a camp-$I(s)$ agent chooses $a=0$ is equal to the benchmark probability $\lambda b_{s}g_{s}/(2g_{s}-1)$, multiplied by the minimal probability that this agent will not be able to counter a criticism from the rival camp; and if all agents always choose $a=1,$ this probability is zero. However, as we show below, the unique dynamically stable equilibrium among all PSPE is the one with the least amount of offensive speech.
The proof of Proposition 2 establishes that a PSPE induces a sequence $x=(x_{m})_{m=1}^{n}$ — where $x_{m}$ is the probability that a camp-$i$ agent chooses $a=0$ in state $s_{i}^{m}$ — that satisfies $x_{m}=c_{m}\phi_{m}(x_{m})$ for every m=1,,nm=1,...,n, where $c_{m}$ and $\phi_{m}$ are given by (8). Note that if there is a strictly positive solution to the equation for $m$, then it also solves the equation $x_{m}=\phi_{m}(c_{m}x_{m}).$ This follows from the observation that when $x_{m}>0$, both equations reduce to being linear in $x_{m}.$ We therefore have the following corollary.
Corollary 1
A sequence of probabilities $(x_{m})_{m=1}^{n}$ is consistent with PSPE if for each $m,$ $x_{m}$ solves $x_{m}=\phi_{m}(c_{m}x_{m}).$
This corollary means that we can regard the R.H.S. of the equation $x_{m}=\phi_{m}(c_{m}x_{m})$ as a self-mapping from $x$ to itself, analogous to a best-reply correspondence in standard game theory. The $x_{m}$ on the R.H.S. represents the fraction of agents on either camp who play $a=0$ in the states $s_{1}^{m}$ and $s_{2}^{m}$. The $x_{m}$ on the L.H.S. is induced by the individual agent’s cutoff $v_{m}^{\ast}$, which is his best-response to the population-level behavior. In light of this interpretation of the equilibrium equations, we propose the following notion of dynamic stability.
@@
Put differently, a PSPE is dynamically stable if at every state, a small perturbation in the fraction of agents (in both camps) who choose $a=0$ in any state leads to a smaller change in the probability that an individual agent chooses that action. [^3] Note that $\phi\left(cx\right)$ is monotone in $x$, so when $x$ changes, $\phi\left(cx\right)$ changes in the same direction.

### 16 Mar 2026 v2

@@ 1 Introduction
Whataboutism is a tactic that people sometimes use to deflect a criticism against violations of social norms. Merriam-Webster dictionary defines whataboutism as “the act or practice of responding to an accusation of wrongdoing by claiming that an offense committed by another is similar or worse.” For example, when a prominent public figure from camp A is reprimanded by a member of camp B for mocking the spouse of a camp B politician (or for keeping silent when another prominent member of camp A did so), a whataboutism response can take the following form: “Why are you criticizing me? Remember when a prominent member of your party ridiculed the child of a politician from my camp? I don’t recall seeing anybody from your party condemning her.”
Whatboutism is a modern-day instance of an ancient rhetorical device known as tu quoque — which Merriam-Webster dictionary defines as “a retort charging an adversary with being or doing what the adversary criticizes in others.” There is a popular sentiment that whataboutism has become increasingly common in recent decades (e.g., see Dykstra (2020)), and that it has sullied public discourse and eroded the liberal ethos that we can agree over means even when we disagree over ends. In Sucio (2025), the Forbes journalist Peter Sucio wrote: [^1] https://www.forbes.com/sites/petersuciu/2025/04/04/whataboutism-has-made-civil-debate-on-social-media-nearly-impossible/ “No matter the topic – especially those about politics and world affairs – it is impossible to have any civil debate on social media without someone responding by making a counter-accusation. This isn’t new, but it has taken such a deep foothold in American politics in recent years, and it isn’t hard to see why.” Sucio went on to attribute the spread of whataboutism to the rise of social media and political polarization.
We propose a formalization of whataboutism as an equilibrium phenomenon, and use this formalism to explore the driving forces behind the phenomenon and its implications for public discourse. In constructing our model, we are guided by three desiderata. First, context should matter: Insulting a soldier on Memorial Day is more offensive than on normal days, and whataboutism arguments would make use of this distinction. Second, the model should enable criticism to be directed not only at offensive speech, but also at subsequent failures to condemn it from within the offender’s camp. Finally, availability of the whataboutism argument should be endogenous: The scenario referred to by one camp’s whataboutism argument is drawn from past speech acts performed by agents on the other camp, which in turn obeyed the same norms shaped by whataboutism. This means that in our model, the social cost of offensive speech should be a function of agents’ equilibrium beliefs regarding the behavior of others.
@@
A condemnation’s effectiveness depends on whether it is internal or external. Condemnation by a member of one’s own camp is always effective; it automatically inflicts a cost on an agent who performed/supported offensive speech. In contrast, external condemnation can be rebuffed with a whataboutism argument. The argument is based on drawing a play path from the distribution induced by equilibrium strategies, in a state that satisfies three requirements: $(1)$ the state is at least as sensitive as the current one; $(2)$ the initial speech-act choice is faced by an agent from the rival camp; and $(3)$ that agent performs an offensive speech act and meets no internal rebuke along the play path. This definition captures whataboutism arguments, which take the following general form: “In a different context that is at least as sensitive as the present one, someone from your camp performed an offensive speech act and no one from your camp condemned it.” The agent’s cost from external condemnation is defined as the probability that he cannot deflect it with whataboutism. The fact that the cost is determined by the ex-ante equilibrium distribution over play paths is what turns our model into a psychological game.
@@ 2 A Model
There are two rival camps, $1$ and $2$, each consisting of a measure one of agents. We use i-i to denote the rival of camp $i$. There is a set of states $S=S_{1}\cup S_{2}$, where $S_{i}=\left\{s_{i}^{1},...,s_{i}^{n}\right\}$ and $S_{1}\cap S_{2}=\emptyset$. The superscript m=1,,nm=1,...,n represents the degree of sensitivity to offensive speech that is associated with the state. When $m^{\prime}\geq m$, we sometimes write $s_{i}^{m^{\prime}}\succsim s_{j}^{m}$ for any $i,j$, and say that $s_{i}^{m^{\prime}}$ is at least as sensitive as $s_{j}^{m}$.
Denote $I(s)=i$ when $s\in S_{i}$. In state $s$, camp $I(s)$ gets an opportunity to perform an offensive speech act targeted at the other camp. In state $s_{i}^{m}$, the act gives intrinsic utility $v\sim U[0,g_{m}]$ to a member of camp $i$, and intrinsic disutility drawn from $U[0,b_{m}]$ to a member of camp i-i. Assume $1<g_{n}<\cdots<g_{1}<2$ and $1<b_{1}<\cdots<b_{n}<2$. Thus, offensive speech in a more sensitive state generates lower intrinsic utility for offenders and higher disutility for victims.
@@
Agents from camp I(s)-I(s) follow an exogenous rule (which we endogenize below): Condemning camp $I(s)$ with probability $\lambda b_{s}$, where $\lambda\in(0,\frac{1}{2})$. The only endogenous objects are the strategies of camp $I(s)$ agents. Since agents from the rival camp I(s)-I(s) are selected at each stage with probability $\frac{1}{2}$ and terminate the game with probability $\lambda b_{s}$, it follows that any profile $\sigma$ of strategies by camp $I(s)$ agents induces a well-defined probability distribution over finite terminal histories. For every $s\in S$, let $\alpha_{\sigma}(s)>0$ denote the ex-ante probability that the agent who terminates the game in $s$ is from camp I(s)-I(s), under the strategy profile $\sigma$.
@@
  • Finally (this is where the psychological-game component enters), when the agent plays $a=1$ and his immediate successor is from the rival camp I(s)-I(s) and plays $a=0$, the agent’s payoff is defined relative to a fixed strategy profile $\sigma$. Specifically, his payoff is $v$ if he can mount a “rebuttal”, and v1v-1 if he cannot. A rebuttal consists of a selection of a state $s^{\prime}\succsim s$ for which $I(s^{\prime})=-I(s)$, and a terminal history in that state — drawn randomly from the distribution over terminal histories induced by $\sigma$ — such that according to $\sigma,$ the stage-$1$ agent in that history plays $a=1$, and all subsequent agents from camp $I(s^{\prime})$ play $a=1$ in that terminal history.
@@
The interpretation of the psychological-game component is as follows. When an agent from camp I(s)-I(s) criticizes an agent from camp $I(s)$ for performing/supporting an offensive speech act, the latter agent can employ a “whataboutism” counterargument: In circumstances that are at least as sensitive as $s$, an agent from the critic’s own camp engaged in offensive speech and faced no subsequent internal rebuke. The psychological-game component arises because the rebuttal is based on a random draw from the distribution over play paths induced by the strategy profile $\sigma$. The rebutting agent needs to retrieve a concrete scenario that would arm him with the whataboutism argument. In this scenario, the state needs to be at least as sensitive as $s$; an agent from the rival camp I(s)-I(s) needs to perform an offensive speech act; and in the ensuing play path, there needs to be no internal condemnation, such that the agent who terminates the game is from camp $I(s)$.
Our modeling of the agent’s retrieval process is based on two desiderata: We want to assume that the agent has some control over the retrieval, but also that success is uncertain. Consequently, our model mixes elements of “directed” and “random” search, in analogy to the macro-theory literature (see Wright et al. (2021)). The agent deliberately targets a specific state, but then he — or someone from his own camp who argues on his behalf — randomly retrieves from memory a concrete state-specific scenario. Thus, when the agent plays $a=1$, he does not know for sure whether he will be able to successfully use whataboutism to rebuff out-group condemnation, but his choice of which state to target is optimal in this regard.
@@
Why do we interpret $a$ as a speech act, rather than as a physical action? Imagine that the whataboutism rebuttal were infeasible. In this case, we would be able to interpret $a=1$ as a physical action that benefits one’s own camp and harms the rival camp. An agent who takes this action incurs a cost if his immediate successor plays $a=0$. This is a fairly standard “network externality”: The agent’s benefit from $a=1$ depends on whether other agents take this action, too.
@@ 3 A Benchmark
There is a unique Subgame Perfect Equilibrium in the benchmark model. In each state $s$, each camp $I(s)$ agent plays $a=1$ if and only if his intrinsic payoff $v$ is above the cutoff
@@
Suppose $v^{k}=0$. This means that camp $I(s)$ agents with $v<\frac{1}{2}\lambda b$ play $a=1$. Such agents earn a negative payoff, regardless of $v^{k+1}$, a contradiction. Now suppose $v^{k}=g$. This means that camp $I(s)$ agents with $v>1$ play $a=0$, even though playing $a=1$ would generate a strictly positive payoff, a contradiction. $\square$
@@
The equilibrium fraction of camp $I(s)$ agents who do not perform/support offensive speech in state $s$ is
@@
This fraction is always in $(0,1)$. The fact that $x_{s}>0$ for every $s$ means that norms of public discourse never break down entirely — i.e., there is always a positive fraction of agents who refrain from offensive speech or criticize it internally. Note that $x(s)$ is increasing in the parameter $b_{s}$ that determines the disutility from offensive speech that members of camp I(s)-I(s) incur. It is also increasing in the parameter $\lambda$ that affects camp $I(s)$ agents’ exposure to external criticism. Finally, it is decreasing in the parameter $g_{s}$ that determines the intrinsic utility that camp $I(s)$ agents derive from verbally attacking the rival camp.
@@ 4 Analysis: Whataboutism
This section characterizes equilibrium behavior in the main model. As a preliminary step, we derive an expression for agents’ psychological expected payoffs. Fix a strategy profile $\sigma$ and consider a state $s$. Consider a camp $I(s)$ agent with valuation $v$, who is selected to move in some stage $k$. If the agent chooses $a=0$ and thus terminates the game, his payoff is by definition 0. Suppose now that the agent chooses $a=1$. Let $\beta_{\sigma}$ denote the probability that the stage-$(k+1)$ agent will choose $a=0$, given the history, conditional on being from the same camp $I(s)$. Recall that $\alpha_{\sigma}(s^{\prime})$ is the ex-ante probability (induced by $\sigma$) that the agent who terminates the game in $s^{\prime}$ is from camp $I(s^{\prime})$. This is the probability that a randomly sampled state-$s^{\prime}$ play path will support a whataboutism rebuttal by a member of camp $I(s^{\prime})$. Then, the stage-$k$ agent’s psychological expected payoff given $\sigma$ from playing $a=1$ is
@@
The first term in this expression is the agent’s intrinsic payoff from performing/supporting the speech act. The second term represents the probability that the agent faces condemnation from within his own camp, which by assumption is impossible to deflect. The third term represents the probability that the agent faces external condemnation which he cannot rebut with a whataboutism argument. Note that the minimum operator is due to the agent’s ability to select among the eligible states (i.e., those that are at least as sensitive as $s$) to maximize the probability of being able to say: “But what about state $s^{\prime}$? An agent from your camp performed an offensive speech act in that state, which is at least as sensitive as $s$, and I didn’t see anyone from your camp criticizing him.”
@@
Let $S=\{s_{1},s_{2}\}$ where $s_{1}\sim s_{2}$, such that both states are associated with the same $(g,b)$. We focus on stationary and symmetric PSPE — i.e., there is a cutoff $v^{\ast}$ such that at every $s$, a camp $I(s)$ agent with valuation $v$ chooses $a=1$ at any history, if and only if $v\geq v^{\ast}$. In such an equilibrium, $\beta$ (the probability that a camp $I(s)$ agent plays $a=0$) is given by $v^{\ast}/g$, independently of the history. Therefore,
@@
The fraction of camp $I(s)$ agents who play $a=0$ in state $s$ at any stage is $x^{\ast}=v^{\ast}/g$. Note that since $g>1$ and $b\lambda<1$, this fraction $x^{\ast}$ is equal to the benchmark level (4) multiplied by a “discount factor” $b\lambda-2g+2<1$. This factor captures an equilibrium effect of whataboutism, which raises the fraction of agents who perform/support offensive speech acts. When $b\lambda-2g+2<0$ there is no interior solution, and $x^{\ast}=0$ in the unique equilibrium — i.e., all norms that constrain public expression break down.
@@ 4.1 The Main Result
For every $m^{\ast}\in\{M,...,n+1\}$, there is a PSPE such that: (i) for every $m<m^{\ast}$, camp $I(s_{i}^{m})$ agents always play $a=1$ in state $s_{i}^{m}$; (ii) for every $m\geq m^{\ast}$, a camp $I(s_{i}^{m})$ agent with valuation $v$ plays $a=1$ if and only if $v\geq v_{m}^{\ast}=c_{m}\theta_{m}g_{m}$; and (iii) the whataboutism rebuttal in any state $s_{i}^{m}$ selects the state $s_{-i}^{m}$. There are no other equilibria.
@@
By essentially the same arguments as in Step 1 of the proof of Proposition 1, for very state $s$ and every stage $k$, there is a cutoff $v_{s}^{k}$ such that a camp $I(s)$ agent with valuation $v$ who moves in stage $k$ plays $a=1$ if and only if $v\geq v_{s}^{k}$.
Given a PSPE $\sigma,$ $\mu_{s}\left(\sigma\right):=\min_{s^{\prime}\succsim s\mid I(s^{\prime})=-I(s)}(1-\alpha_{\sigma}(s^{\prime}))$ is the probability that external condemnation of camp $I(s)$ in state $s$ cannot be rebuffed by whataboutism. Then, in state $s$, a camp $I(s)$ agent with valuation $v$ who moves in stage $k$ earns a payoff of
@@
By Step 1, in equilibrium, every state $s$ is associated with a stationary cutoff $v_{s}^{\ast}$ that governs camp $I(s)$ agents’ behavior. Therefore, for every state $s^{\prime}$ where $I(s^{\prime})\neq I(s)$, we obtain essentially the same expression as (6), repeated here for convenience:
@@
It is now straightforward to confirm that the sequences $(x_{m})$ and $(y_{m})$ described by Step 4 solve (11). Therefore, Step 4 fully characterizes the set of PSPE. It also means that for any state $s_{i}^{m}$, the state $s^{\prime}$ that minimizes $1-\alpha_{\sigma}(s^{\prime})$ among eligible ones is $s_{-i}^{m}$ — i.e., a whataboutism rebuttal always targets an equivalent-sensitivity state. $\blacksquare$
Proposition 2 means that any PSPE is characterized by a “sensitivity threshold” $m^{\ast}\geq M$. In state $s_{i}^{m},$ $m\geq m^{\ast}$, the probability that a camp $i$ agent refrains from offensive speech (and condemns his comrades when they do not) is $x_{m}=v_{m}^{\ast}/g_{m}$, which is equal to the corresponding probability in the benchmark ($c_{m}$) multiplied by the “discount factor” $\theta_{m}.$ Since both $c_{m}$ and $\theta_{m}$ increase with $m,$ the probability of condemnation increases with the sensitivity of the state. The term $\theta_{m}$ can be interpreted as the “whataboutism effect”, which lowers the likelihood that agents condemn members of their camp for performing/supporting offensive speech. The whataboutism effect can be dramatic and lead to a “breakdown of norms” in the sense that in if $\theta_{m}<0$ for some state $s_{i}^{m},$ then in all weakly less sensitive states, all agents (regardless of their camp) perform/support offensive speech.
@@
The PSPE are ranked by their amount of offensive speech: The higher $m^{\ast}$, the higher that amount. In particular, there is always an equilibrium with complete breakdown of norms, where $m^{\ast}=n+1$ such that agents always engage in (and never condemn) offensive speech. To see why, note that the equilibrium probability that in state $s$ a camp $I(s)$ agent chooses $a=0$ is equal to the benchmark probability $\lambda b_{s}g_{s}/(2g_{s}-1)$, multiplied by the minimal probability that this agent will not be able to counter a criticism from the rival camp; and if all agents always choose $a=1,$ this probability is zero. However, as we show below, the unique dynamically stable equilibrium among all PSPE is the one with the least amount of offensive speech.
The proof of Proposition 2 establishes that a PSPE induces a sequence $x=(x_{m})_{m=1}^{n}$ — where $x_{m}$ is the probability that a camp $i$ agent chooses $a=0$ in state $s_{i}^{m}$ — that satisfies $x_{m}=c_{m}\phi_{m}(x_{m})$ for every m=1,,nm=1,...,n, where $c_{m}$ and $\phi_{m}$ are given by (8). We can regard the R.H.S. of the equation $x_{m}=c_{m}\phi_{m}(x_{m})$ as a self-mapping from $x$ to itself, analogous to a best-reply correspondence in standard game theory. The $x_{m}$ on the R.H.S. represents the fraction of agents on either camp who play $a=0$ in the states $s_{1}^{m}$ and $s_{2}^{m}$. The $x_{m}$ on the L.H.S. is induced by the individual agent’s cutoff $v_{m}^{\ast}$, which is his best-response to the population-level behavior. In light of this interpretation of the equilibrium equations, we propose the following notion of dynamic stability.
@@
Put differently, a PSPE is dynamically stable if at every state, a small perturbation in the fraction of agents (in both camps) who choose $a=0$ in any state leads to a smaller change in the probability that an individual agent chooses that action. [^3] Note that $\phi\left(x\right)$ is monotone in $x$, so when $x$ changes, $\phi\left(x\right)$ changes in the same direction.